OpenStudy (anonymous):

If 1200 cm squared of material is available to make a box with a square base and an open top, find the largest possible volume of the box. thank you

OpenStudy (radar):

The box will be formed by cutting in each side at the corners for a distance x and folding up the sides to form the box. The depth of the box will be x, the length and width will be 1200-2x. Units are cm. Volume V=x(1200-2x)(1200-2x) V=x)1440000-4800x+4x^2)

OpenStudy (radar):

\[V=4x ^{3}-4800x ^{2}+1440000\]

OpenStudy (radar):

Differentiate and set to for max min

OpenStudy (radar):

\[12x ^{2}-9600x=0\]

OpenStudy (radar):

Solve for x the depth, the sides would be 1200-2x. Now if the material is 1200 sq cm and not a piece that is 1200 x1200 then use sides of 1200^sqrt.

OpenStudy (radar):

The way I read it, was a piece of material 1200 by 1200

OpenStudy (radar):

I see where I made a mistake in the multiplying failing to multiplying byx

OpenStudy (anonymous):

thank you thank you! In this question, I'm given the surface area --> 1200 = w^2 + 4 ( h x w )

OpenStudy (anonymous):

I apologize if I made the question confusing =(

OpenStudy (radar):

O.K. I'm sorry. does the problem say you have a square piece of material that is 1200 cm sq?

OpenStudy (anonymous):

Yes that is correct. I'm sorry for the confusion

OpenStudy (radar):

Or is the completed box to have 1200 sq cm of surface area

OpenStudy (anonymous):

Hm... let me read again. I'm sorrry!!!

OpenStudy (anonymous):

Hm... let me read again. I'm sorrry!!!

OpenStudy (anonymous):

if \[1200 cm ^{2 }\] of material is available to make a box with a square base and an open top, find the largest possible volume of the box

OpenStudy (radar):

I am sorry, but I have to go pick up my car. If the problem is not solved when I get back tonight I will make furthert attempts to help. Thanks for clarify. I think I know what you are to do.

OpenStudy (anonymous):

Thank you so much for you help sir!

OpenStudy (radar):

I'm back. if you have 1200 cm^2 of material then you have a square whose sides are equal to the square root of 1200 or 34.64 cm. The box will be formed by cutting in each side at the corners for a distance x and folding up the sides to form the box. The depth of the box will be x, the length and width will be 34.64-2x. Units are cm.

OpenStudy (radar):

Volume V=x(34.64-2x)(34.64-2x) V=(x)(1200-138.56x+4x^2) V=1200x-138.5x^2+4x^3

OpenStudy (radar):

\[V=4x ^{3}-138.5x ^{2}+1200x\]

OpenStudy (radar):

Now differentiate and set to zero

OpenStudy (radar):

\[12x ^{2}-277x+1200=0\]

OpenStudy (radar):

Now you can use quadratic equation and solve for x which represents the depth and the sides would be 34.64-2x. Good luck